OpenStudy (andpeggy):

A manufacturing plant is going to use two different stamping machines to complete an order of 975 units. One produces 100 units per hour, while the other produces 75 units per hour. How long must each machine operate to complete the order if ,during the process, the faster machine is shut down for 25 hours for repairs? ( I just need help setting it up, thank you!)

OpenStudy (badsheep):

let t = time operated by the slower machine then (t-2.5)= time operated by the faster machine : 75t + 100(t-2.5) = 975 75t + 100t - 250 = 975

OpenStudy (badsheep):

175t = 975 + 250 175t = 1225 t = 1225/175 = t = 7 hrs by the slower machine and 7 - 2.5 = 4.5 hrs by the faster

OpenStudy (andpeggy):

thanks!